A company begins a review of ordering policies for its continuous review system by checking the current policies for a sample of SKUs. Following are the characteristics of one​ item:

follows≻Demand

​(D) =

80 units/week (Assume

50 weeks per​ year)

follows≻Ordering

and setup cost​ (S) =90​/order

follows≻Holding

cost​ (H) =18​/unit/year

follows≻Lead

time​ (L) =

22

​week(s)

follows≻Standard

deviation of weekly demand​ =

2020

units

follows≻​Cycle-service

level​ =

9292

percent

follows≻EOQ

​=

200200

units

Using the above​ information, develop the best policies for a periodic review system. Refer to the standard normal table

LOADING...

for​ z-values.

The table below shows the total area under the normal curve for a point that is Z standard deviations to the right of the mean.

Z

0.00

0.01

0.02

0.03

0.04

0.05

0.06

0.07

0.08

0.09

0.0

0.5000

0.5040

0.5080

0.5120

0.5160

0.5199

0.5239

0.5279

0.5319

0.5359

0.1

0.5398

0.5438

0.5478

0.5517

0.5557

0.5596

0.5636

0.5675

0.5714

0.5754

0.2

0.5793

0.5832

0.5871

0.5910

0.5948

0.5987

0.6026

0.6064

0.6103

0.6141

0.3

0.6179

0.6217

0.6255

0.6293

0.6331

0.6368

0.6406

0.6443

0.6480

0.6517

0.4

0.6554

0.6591

0.6628

0.6664

0.6700

0.6736

0.6772

0.6808

0.6844

0.6879

0.5

0.6915

0.6950

0.6985

0.7019

0.7054

0.7088

0.7123

0.7157

0.7190

0.7224

0.6

0.7258

0.7291

0.7324

0.7357

0.7389

0.7422

0.7454

0.7486

0.7518

0.7549

0.7

0.7580

0.7612

0.7642

0.7673

0.7704

0.7734

0.7764

0.7794

0.7823

0.7852

0.8

0.7881

0.7910

0.7939

0.7967

0.7996

0.8023

0.8051

0.8079

0.8106

0.8133

0.9

0.8159

0.8186

0.8212

0.8238

0.8264

0.8289

0.8315

0.8340

0.8365

0.8389

1.0

0.8413

0.8438

0.8461

0.8485

0.8508

0.8531

0.8554

0.8577

0.8599

0.8621

1.1

0.8643

0.8665

0.8686

0.8708

0.8729

0.8749

0.8770

0.8790

0.8810

0.8830

1.2

0.8849

0.8869

0.8888

0.8907

0.8925

0.8944

0.8962

0.8980

0.8997

0.9015

1.3

0.9032

0.9049

0.9066

0.9082

0.9099

0.9115

0.9131

0.9147

0.9162

0.9177

1.4

0.9192

0.9207

0.9222

0.9236

0.9251

0.9265

0.9279

0.9292

0.9306

0.9319

1.5

0.9332

0.9345

0.9357

0.9370

0.9382

0.9394

0.9406

0.9418

0.9430

0.9441

1.6

0.9452

0.9463

0.9474

0.9485

0.9495

0.9505

0.9515

0.9525

0.9535

0.9545

1.7

0.9554

0.9564

0.9573

0.9582

0.9591

0.9599

0.9608

0.9616

0.9625

0.9633

1.8

0.9641

0.9649

0.9656

0.9664

0.9671

0.9678

0.9686

0.9693

0.9700

0.9706

1.9

0.9713

0.9719

0.9726

0.9732

0.9738

0.9744

0.9750

0.9756

0.9762

0.9767

2.0

0.9773

0.9778

0.9783

0.9788

0.9793

0.9798

0.9803

0.9808

0.9812

0.9817

2.1

0.9821

0.9826

0.9830

0.9834

0.9838

0.9842

0.9846

0.9850

0.9854

0.9857

2.2

0.9861

0.9865

0.9868

0.9871

0.9875

0.9878

0.9881

0.9884

0.9887

0.9890

2.3

0.9893

0.9896

0.9898

0.9901

0.9904

0.9906

0.9909

0.9911

0.9913

0.9916

2.4

0.9918

0.9920

0.9922

0.9925

0.9927

0.9929

0.9931

0.9932

0.9934

0.9936

2.5

0.9938

0.9940

0.9941

0.9943

0.9945

0.9946

0.9948

0.9949

0.9951

0.9952

2.6

0.9953

0.9955

0.9956

0.9957

0.9959

0.9960

0.9961

0.9962

0.9963

0.9964

2.7

0.9965

0.9966

0.9967

0.9968

0.9969

0.9970

0.9971

0.9972

0.9973

0.9974

2.8

0.9974

0.9975

0.9976

0.9977

0.9977

0.9978

0.9979

0.9980

0.9980

0.9981

2.9

0.9981

0.9982

0.9983

0.9983

0.9984

0.9984

0.9985

0.9985

0.9986

0.9986

3.0

0.9987

0.9987

0.9987

0.9988

0.9988

0.9989

0.9989

0.9989

0.9990

0.9990

3.1

0.9990

0.9991

0.9991

0.9991

0.9992

0.9992

0.9992

0.9992

0.9993

0.9993

3.2

0.9993

0.9993

0.9994

0.9994

0.9994

0.9994

0.9994

0.9995

0.9995

0.9995

3.3

0.9995

0.9995

0.9995

0.9995

0.9996

0.9996

0.9996

0.9996

0.9996

0.9997

a. What value of P gives the same approximate number of orders per year as the​ EOQ?

nothing

weeks. ​(Enter your response rounded to the nearest whole​ number.)

b. To provide a 92 percent​ cycle-service level, the average safety stock is

nothing

units. ​(Enter your response rounded to the nearest whole​ number.)

To provide a

92 percent​ cycle-service level, the target inventory level is

nothing

units. ​(Enter your response rounded to the nearest whole​ number.)
Answer #1

EOQ = 200 units

Number of orders per year = annual demand / EOQ = 80x50/200 =20

Time between the orderes = No. of weeks / No of orders = 50/20 =2.5

which is equivalent to P.

Safety stock = z x sigma ( P+L)1/2

= 1.41 x 20 ( 2.5+2)1/2 = 59.82 = 60 units

Order up to level or target inventory level = (P+L) x weekly demand + Safety stock

= 4.5 x80+60 = 420